What is the answer to 2 1/2(x-1)=-95/9?
Answer by rfer(16322)

2.5x-2.5=-95/9
2.5x=-95/9+25/10
25/10x=-950/90+225/90
25/10x=-725/90
x=-725/90*10/25
x=-7250/2250
x=-3 2/9